# Identifying noise transients in gravitational-wave data arising from nonlinear couplings

### v1

@@ Abstract
Noise in various interferometer systems can sometimes couple non-linearly to create excess noise in the gravitational wave (GW) strain data. Third-order statistics, such as bicoherence and biphase, can identify these couplings and help discriminate those occurrences from astrophysical GW signals. However, the conventional analysis can yield large bicoherence values even when no phase-coupling is present, thereby, resulting in false identifications. Introducing artificial phase randomization in computing the bicoherence reduces such occurrences with negligible impact on its effectiveness for detecting true phase-coupled disturbances. We demonstrate this property with simulated disturbances – focusing only on short-duration ones (lasting up to a few seconds) in this work. Statistical hypothesis testing is used for distinguishing phase-coupled disturbances from non-phase coupled ones when employing the phase-randomized bicoherence. We also obtain an expression for the bicoherence value that minimizes the sum of the probabilities of false positives and false negatives. This can be chosen as a threshold for shortlisting bicoherence triggers for further scrutiny for the presence of non-linear coupling. Finally, the utility of the phase-randomized bicoherence analysis in GW time-series data is demonstrated for the following three scenarios: (1) Finding third-order statistical similarities within categories of noise transients, such as blips and koi fish. If these non-Gaussian noise transients, or glitches, have a common source, their bicoherence maps can have similarities arising from common bifrequencies related to that source. (2) Differentiating linear or non-linear phase-coupled glitches from compact binary coalescence signals through their bicoherence maps. This is explained with a simulated signal. (3) Identifying repeated bifrequencies in the second and third observation runs (i.e., O2 and O3) of LIGO and Virgo.
Keywords: Noise-artifacts, Non-linear Couplings, Bispectrum, Bicoherence, Binary Black Holes
@@ 1 Introduction
The detectors of the Laser Interferometer Gravitational-wave Observatory (LIGO) [1], Virgo [2] and the Kamioka Gravitational Wave Detector (KAGRA) [3] are among the most sophisticated scientific instruments in the world. To date, LIGO and Virgo have detected several dozens of gravitational wave (GW) signals [4, 5, 6, 7] from colliding binaries involving black holes and neutron stars. These detectors on the globe are sensitive to GW signals in the $10-10^{4}$ Hz band [8], but are often plagued by unwanted noise transients from various terrestrial sources, which can adversely impact that sensitivity [9, 10, 11, 12]. The quality of data or search sensitivity can be improved by removing or minimizing the noise transients in the instruments, whenever possible. It is, however, non-trivial to do so because the origins of many of them are unknown or, in some cases, not in our control (e.g., earthquakes) [13]. Those difficulties notwithstanding, detector characterization exercises and related data-analysis techniques have been employed for mitigating several kinds of noise transients in GW searches [14, 15, 16]. In some of these cases the excess noise-power in the GW strain-data is due to bilinear or nonlinear couplings among various detector subsystems and components [17, 18, 19, 20, 21, 22]. Whenever the causes of disturbances are found, steps are taken to remove or reduce their impact on the GW strain channel and the detector’s sensitivity to astrophysical signals.
@@
In this work, we develop a technique for detecting the presence of nonlinear couplings at second order, i.e., quadratic nonlinearities, in the GW strain data. A third-order statistic that is useful for such identification is the bicoherence [18]. The traditional bicoherence [23] construction, however, has certain limitations that can often result in misidentifcation of phase-coupled noise disturbances. To rectify it, we introduce the phase-randomized (PR) bicoherence analysis to GW data, and demonstrate its utility with simulated signal studies in Sec. 2. In practice, random noise can sometimes mimic the bicoherence response of phase-coupled (PC) disturbances, thereby leading to false detections, or false alarms, even when phase randomization is used. To determine whether there is evidence of phase-coupling in the data or not, a hypothesis test is carried out between PC and non-phase-coupled (NPC) disturbances by employing the PR bicoherence analysis. Along the way, we derive an expression for the bicoherence value that minimizes the false identification of non-phase-coupled disturbances as phase-coupled ones in Sec. 3. This is useful for assessing how significant any phase-coupled disturbance is based on its bicoherence value. In Sec. 4, we apply the PR bicoherence analysis to GW data from LIGO detectors, and the utility of this analysis to identify nonlinear couplings is explained in terms of three examples in real data.
@@ 2 Higher order statistics -- Quadratic Phase Coupling
where $f_{k}$, $f_{l}$, $f_{m}$ are taken to be discrete frequencies, with $k$ and $l=1,...,N$, $k\neq l$, and $f_{m}=f_{k}\pm f_{l}$, such that $m\leq N$. Here, $N$ is the Nyquist frequency index and $f_{N}$ will denote the Nyquist frequency. Thus, the above frequencies obey the properties: (i) $0\leq f_{k,l}$ $\leq f_{N}$; (ii) $0\leq f_{k}\pm f_{l}\leq f_{N}$. The bispectrum is computed for pairs of frequencies $(f_{k},f_{l})$ – also called bifrequencies – and obeys the following properties: (i) $B(f_{k},f_{l})=B^{*}(-f_{k},-f_{l})=B(f_{l},f_{k})$; and (ii) $B(f_{k},f_{l})=B(f_{k},-f_{k}-f_{l})$ [25]. If $\tilde{Z}$ has a signal that is phase-coupled with $\tilde{X}(f_{k})$ and $\tilde{Y}(f_{l})$, then the bispectrum can assume large values at the bifrequency $(f_{k},f_{l})$ if that coupling is strong. A normalized version of the bispectrum is defined as:
@@
and is called the bicoherence [26]. Here, $|\tilde{A}(f_{l})|^{2}\equiv\tilde{A}_{\rm\alpha}(f_{l})\tilde{A}^{*}_{\rm\alpha}(f_{l})$ and $E[\tilde{A}(f_{l})]\equiv\frac{1}{M}~{}\sum_{\rm\alpha=1}^{M}~{}\tilde{A}_{\rm\alpha}(f_{l})$. Below, we will most often employ the bicoherence instead of the bispectrum.
@@
In the context of GW data analysis, $x(t)$, $y(t)$ and $z(t)$ are time-series data from the GW strain channel or any of the auxiliary GW detector channels (e.g., associated with instruments and sensors). When one sets $x(t)=y(t)=z(t)$ in Eq. (3), the resulting statistic is called the auto--bicoherence. Otherwise the statistic is also called the cross--bicoherence. In this work, we focus on the auto--bicoherence and briefly present a feasiblity study on the cross--bicoherence in Appendix C.
@@
The bispectrum is manifestly complex and its phase is called the biphase, $\Phi_{\rm d}$. If $\Phi(f_{k})$, $\Phi(f_{l})$, and $\Phi(f_{m})$ denote the phases of $\tilde{X}_{\rm\alpha}(f_{k})$, $\tilde{Y}_{\rm\alpha}(f_{l})$ and $\tilde{Z}_{\rm\alpha}(f_{m})$, respectively, then $\Phi_{\rm d}\equiv\Phi(f_{k})+\Phi(f_{l})-\Phi(f_{m})$. The utility of the variance of the biphase will be discussed in the following section, and is denoted as $\sigma^{\prime~{}2}_{\rm~{}\Phi}$. It is empirically related to the bicoherence [27] as
$\sigma^{\prime~{}2}_{\rm~{}\Phi}\approx\frac{1}{2M}~{}\left[\frac{1}{b^{2}(f_{k},f_{l})}-1\right]\,.$ (4)
Figure 1: The impact of phase randomization on distinguishing nonlinearly phase-coupled signals (labeled “PC” and shown as green dots) from those without (labeled “NPC” and shown as red dots), in plots of biphase variance vs bicoherence. We compute these statistics for the $(20,80)$ Hz bifrequency of the simulated PC and NPC signals described by Eq. (2.1). Each dot represents a different realization of the noise to which those signals are added. Left: Here the bicoherence is computed traditionally, i.e., with no phase randomization (NPR). Many PC signals show up with strong bicoherence values in this plot, but so do many NPC signals. The assumption that the bispectrum phases of the $M$ segments are independent of one another is invalid for these NPC signals and leads to their high bicoherence values, thus, making it difficult to distinguish them from the PC signals. Middle: The bicoherence values of the of PC signals shown in the left figure are minimally affected by the introduction of artificial phase randomization (PR). Right: For the same NPC signals shown in the left figure, the use of phase randomization helps in identifying the lack of phase coupling in those cases and gets their bicoherence to stay low. (The corresponding biphase variance tends to rise.) This helps in distinguishing the NPC signals from the PC ones when phase randomization is applied.
@@
Given a bifrequency, if the corresponding biphases of the $M$ segments happen to be independent of one another, and have a uniform distribution in the interval $(-\pi,\pi]$, then its bispectrum $\approx 0$. Its bicoherence will be negligible as well. On the other hand, any bifrequency with a zero biphase will result in a non-vanishing bispectrum, and a bicoherence $\approx 1$.
@@
Unfortunately, the bicoherence – as traditionally defined in Eq. (3) –suffers from the limitation that it can turn out to be unity even when there is no phase coupling. This can happen, e.g., when $\Phi_{\rm d}$ is a constant across all $M$ segments. In that event, the bicoherence is just $|E[e^{i{\Phi_{\rm d}}}]|=|e^{i{\Phi_{\rm d}}}|$, which is unity. This would be a false alarm since that value is indistinguishable from the one expected of a phase-coupled signal, where the biphase is 0 and, consequently, the bicoherence is unity. One way to address this problem is to multiply the phase $\Phi_{\rm d}^{\alpha}$ of the data from each segment with a random variable $R^{(\rm\alpha)}$ such that $R^{\rm(\alpha)}$ is the $\alpha$-th realization of a random variable that follows either a uniform distribution over $(-\pi,\pi]$ or a zero-mean normal distribution with a variance $\sigma^{2}_{\rm R}$. The phase-randomized bispectrum is:
@@
and the phase-randomized bicoherence $b_{\rm pr}(f_{k},f_{l})$ is given by Eq. (3) with $B(f_{k},f_{l})$ replaced by $B_{\rm pr}(f_{k},f_{l})$. The PR prescription works equally well regardless of which of the two distributions is used for $R^{\alpha}$ [25]. The expression for the variance of biphase corresponding to phase-randomized bicoherence is similar to Eq. (4). It is estimated by using $b_{\rm pr}(f_{k},f_{l})$ instead of $b(f_{k},f_{l})$, and the variance of its biphase will be denoted as $\sigma_{\Phi}^{2}$.
@@
Below we show with simulations how the bicoherence statistic can be used to discern the presence of quadratic phase coupling. Thereafter, we present some applications in real GW data. Throughout this paper, we fixed the values of some of the analysis parameters for both simulated signals and GW data, unless stated otherwise. The duration of all data chunks analyzed is chosen to be 4 sec, with a sampling frequency $f_{\rm s}=1024$ Hz; of course, the Nyquist frequency is set to $f_{N}=\frac{f_{\rm s}}{2}$. For computing the bicoherence, each time-series data chunk is divided into $M$ segments and their Fourier transforms taken after applying the Hanning window on them. We also use $98\%$ overlap among consecutive segments.
The optimal choice for the number of segments ($M$) is decided empirically. For a fixed chunk duration, if $M$ is small, then the relative contribution of noise to the bicoherence magnitude can be large. On the other hand, if $M$ is large, then the overlap of the segments will be large for the same chunk duration. In such a case, the different segments will cease to be independent and the bicoherence for a real signal, of a similar duration as the chunk, will plateau. By performing injection of simulated signals in real data and studying known short duration couplings (lasting no more than a few seconds) we found $M\approx 31$.

### v2

@@ Abstract
Noise in various interferometer systems can sometimes couple non-linearly to create excess noise in the gravitational wave (GW) strain data. Third-order statistics, such as bicoherence and biphase, can identify these couplings and help discriminate those occurrences from astrophysical GW signals. However, the conventional analysis can yield large bicoherence values even when no phase-coupling is present, thereby, resulting in false identifications. Introducing artificial phase randomization in computing the bicoherence reduces such occurrences with negligible impact on its effectiveness for detecting true phase-coupled disturbances. We demonstrate this property with simulated disturbances – focusing only on short-duration ones (lasting up to a few seconds) and employing mainly the auto-bicoherence in this work. Statistical hypothesis testing is used for distinguishing phase-coupled disturbances from non-phase coupled ones when employing the phase-randomized bicoherence. We also obtain an expression for the bicoherence value that minimizes the sum of the probabilities of false positives and false negatives. This can be chosen as a threshold for shortlisting bicoherence triggers for further scrutiny for the presence of non-linear coupling. Finally, the utility of the phase-randomized bicoherence analysis in GW time-series data is demonstrated for the following three scenarios: (1) Finding third-order statistical similarities within categories of noise transients, such as blips and koi fish. If these non-Gaussian noise transients, or glitches, have a common source, their bicoherence maps can have similarities arising from common bifrequencies related to that source. (2) Differentiating linear or non-linear phase-coupled glitches from compact binary coalescence signals through their bicoherence maps. This is explained with a simulated signal. (3) Identifying repeated bifrequencies in the second and third observation runs (i.e., O2 and O3) of LIGO and Virgo.
@@ 1 Introduction
The detectors of the Laser Interferometer Gravitational-wave Observatory (LIGO) [1], Virgo [2] and the Kamioka Gravitational Wave Detector (KAGRA) [3] are among the most sophisticated scientific instruments in the world. To date, LIGO and Virgo have detected several dozens of gravitational wave (GW) signals [4, 5, 6, 7] from colliding binaries involving black holes and neutron stars. These detectors on the globe are sensitive to GW signals in the $10-10^{4}$ Hz band [8], but are often plagued by unwanted noise transients from various terrestrial sources, which can adversely impact that sensitivity [9, 10, 11, 12]. The quality of data or search sensitivity can be improved by removing or minimizing the noise transients in the instruments, whenever possible. It is, however, non-trivial to do so because the origins of many of them are unknown or, in some cases, not in our control (e.g., earthquakes) [13]. Those difficulties notwithstanding, detector characterization exercises and related data-analysis techniques have been employed for mitigating several kinds of noise transients in GW searches [14, 15, 16, 17, 18, 19, 20, 21]. In some of these cases the excess noise-power in the GW strain-data is due to bilinear or nonlinear couplings among various detector subsystems and components [22, 23, 24, 25, 26, 27]. These include recent applications of machine learning algorithms such as NonSENS [27] and DeepClean [26], which have been developed to variously regress linear, nonlinear, and nonstationary noise that typically last for tens of seconds or longer. For that purpose, certain “witness” sensor data are used to estimate instrumental or environmental noise contributions to the GW strain channel and regress them. These activities have demonstrated varying degrees of success in lowering the strain noise in software and improving the sensitivity to astrophysical searches.
@@
In this work, we develop a technique for detecting the presence of nonlinear couplings at second order, i.e., quadratic nonlinearities, in the GW strain data. A third-order statistic that is useful for such identification is the bicoherence [23], specifically associated with non-Gaussian transients, or “glitches”. The traditional bicoherence [28] construction, however, has certain limitations that can often result in misidentifcation of phase-coupled noise disturbances. To rectify it, we introduce the phase-randomized (PR) bicoherence analysis to GW data, and demonstrate its utility with simulated signal studies in Sec. 2. In practice, random noise can sometimes mimic the bicoherence response of phase-coupled (PC) disturbances, thereby leading to false detections, or false alarms, even when phase randomization is used. To determine whether there is evidence of phase-coupling in the data or not, a hypothesis test is carried out between PC and non-phase-coupled (NPC) disturbances by employing the PR bicoherence analysis. Along the way, we derive an expression for the bicoherence value that minimizes the false identification of non-phase-coupled disturbances as phase-coupled ones in Sec. 3. This is useful for assessing how significant any phase-coupled disturbance is based on its bicoherence value. In Sec. 4, we apply the PR auto-bicoherence analysis to GW data from LIGO Hanford and Livingston detectors, and the utility of this analysis to identify nonlinear couplings is explained in terms of three examples in real data.
@@ 2 Higher order statistics -- Quadratic Phase Coupling
where $f_{k}$, $f_{l}$, $f_{m}$ are taken to be discrete frequencies, with $k$ and $l=1,...,N$, $k\neq l$, and $f_{m}=f_{k}\pm f_{l}$, such that $m\leq N$. Here, $N$ is the Nyquist frequency index and $f_{N}$ will denote the Nyquist frequency. Thus, the above frequencies obey the following limits: (1) $0\leq f_{k,l}$ $\leq f_{N}$; (2) $0\leq f_{k}\pm f_{l}\leq f_{N}$. The bispectrum is computed for pairs of frequencies $(f_{k},f_{l})$ – also called bifrequencies – and obeys the following properties when $z_{\alpha}(t)$ is chosen to be the same as $y_{\alpha}(t)$, for all $\alpha$ [30]: (1) $B(f_{k},f_{l})=B^{*}(-f_{k},-f_{l})=B(f_{l},f_{k})$; and (2) $B(f_{k},f_{l})=B(f_{k},-f_{k}-f_{l})$. If $\tilde{Z}$ has a signal that is phase-coupled with $\tilde{X}(f_{k})$ and $\tilde{Y}(f_{l})$, then the bispectrum can assume large values at the bifrequency $(f_{k},f_{l})$ if that coupling is strong. A normalized version of the bispectrum is defined as:
@@
and is called the bicoherence [31]. Here, $|\tilde{A}(f_{l})|^{2}\equiv\tilde{A}_{\rm\alpha}(f_{l})\tilde{A}^{*}_{\rm\alpha}(f_{l})$ and $E[\tilde{A}(f_{l})]\equiv\frac{1}{M}~{}\sum_{\rm\alpha=1}^{M}~{}\tilde{A}_{\rm\alpha}(f_{l})$. Below, we will most often employ the bicoherence instead of the bispectrum. It is evident that the bicoherence obeys $0\leq b(f_{k},f_{l})\leq 1$. This is most readily seen by identifying the product $\tilde{X}_{\rm\alpha}\tilde{Y}_{\rm\alpha}$ as the components of a single complex vector, say, $\tilde{W}$. It then follows that
$\frac{B(f_{k},f_{l})}{\sqrt{E\Big{[}\big{|}\tilde{X}(f_{k})~{}\tilde{Y}(f_{l})\big{|}^{2}\Big{]}~{}E\Big{[}\big{|}\tilde{Z}(f_{m})\big{|}^{2}\Big{]}}}=\frac{\sum_{\rm\alpha=1}^{M}~{}\tilde{W}_{\rm\alpha}~{}\tilde{Z}_{\rm\alpha}^{*}}{\sqrt{\sum_{\rm\alpha=1}^{M}~{}\left|\tilde{W}_{\rm\alpha}\right|^{2}}\sqrt{\sum_{\rm\alpha=1}^{M}~{}\left|\tilde{Z}_{\rm\alpha}\right|^{2}}}\,,$
which is just the inner product of two unit-norm complex vectors. The magnitude of such a product obeys the Cauchy-Schwarz inequality, which is just the constraint noted above on $b(f_{k},f_{l})$.
@@
In the context of GW data analysis, $x(t)$, $y(t)$ and $z(t)$ are time-series data from the GW strain channel or any of the auxiliary GW detector channels (e.g., associated with instruments and sensors). When one sets $x(t)=y(t)=z(t)$ in Eq. (3), the resulting statistic is called the auto--bicoherence. Otherwise the statistic is also called the cross--bicoherence.
@@
The bispectrum is manifestly complex and its phase is called the biphase, $\Phi_{\rm d}$. The variance of the biphase is denoted as $\sigma^{\prime~{}2}_{\rm~{}\Phi}$. For the auto-bicoherence, $\Phi_{\rm d}$ is empirically related to it as [32]
$\sigma^{\prime~{}2}_{\rm~{}\Phi}\approx\frac{1}{2M}~{}\left[\frac{1}{b^{2}(f_{k},f_{l})}-1\right]\,,$ (4)
@@
and will be discussed further in the following section.On the other hand, if $\Phi^{\alpha}(f_{k})$, $\Phi^{\alpha}(f_{l})$, and $\Phi^{\alpha}(f_{m})$ denote the phases of $\tilde{X}_{\rm\alpha}(f_{k})$, $\tilde{Y}_{\rm\alpha}(f_{l})$ and $\tilde{Z}_{\rm\alpha}(f_{m})$, respectively, then $\Phi^{\alpha}_{\rm d}\equiv\Phi^{\alpha}(f_{k})+\Phi^{\alpha}(f_{l})-\Phi^{\alpha}(f_{m})$ is the biphase of individual segments. Given a bifrequency, if the corresponding biphases of the $M$ segments happen to be independent of one another, and have a uniform distribution in the interval $(-\pi,\pi]$, then its bispectrum $\approx 0$. Its bicoherence will be negligible as well. On the other hand, when $\Phi_{d}^{\alpha}\approx 0$ for every segment, then it will result in a non-vanishing bispectrum and a bicoherence $\approx 1$, which is just the maximum value allowed by the Cauchy-Schwarz inequality.
Figure 1: The impact of phase randomization on distinguishing nonlinearly phase-coupled signals (labeled “PC” and shown as green dots) from those without (labeled “NPC” and shown as red dots), in plots of biphase variance vs (auto-)bicoherence. We compute these statistics for the $(20,80)$ Hz bifrequency of the simulated PC and NPC signals described by Eq. (2.1). Each dot represents a different realization of the noise to which those signals are added. Left: Here the (auto-)bicoherence is computed traditionally, i.e., with no phase randomization (NPR). Many PC signals show up with strong bicoherence values in this plot, but so do many NPC signals. The assumption that the bispectrum phases of the $M$ segments are independent of one another is invalid for these NPC signals and leads to their high bicoherence values, thus, making it difficult to distinguish them from the PC signals. Middle: The bicoherence values of the of PC signals shown in the left figure are minimally affected by the introduction of artificial phase randomization (PR). Right: For the same NPC signals shown in the left figure, the use of phase randomization helps in identifying the lack of phase coupling in those cases and gets their bicoherence to stay low. (The corresponding biphase variance tends to rise.) This helps in distinguishing the NPC signals from the PC ones when phase randomization is applied.
@@
Unfortunately, the bicoherence – as traditionally defined in Eq. (3) – suffers from the limitation that it can turn out to be unity even when there is no phase coupling. This can happen, e.g., when the $\Phi_{\rm d}^{\alpha}$ are the same across all $M$ segments and, say, equal $\Phi_{\rm d}$. In that event, the bicoherence is just $|E[e^{i{\Phi_{\rm d}}}]|=|e^{i{\Phi_{\rm d}}}|$, which is unity. This would be a false alarm since that value is indistinguishable from the one expected of a phase-coupled signal, where the biphase is 0 and, consequently, the bicoherence is unity. One way to address this problem is to multiply the phase $\Phi_{\rm d}^{\alpha}$ of the data from each segment with a random variable $R^{(\rm\alpha)}$ such that $R^{\rm(\alpha)}$ is the $\alpha$-th realization of a random variable that follows either a uniform distribution over $(-\pi,\pi]$ or a zero-mean normal distribution with a variance $\sigma^{2}_{\rm R}$. The phase-randomized bispectrum is:
@@
and the phase-randomized bicoherence $b_{\rm pr}(f_{k},f_{l})$ is given by Eq. (3) with $B(f_{k},f_{l})$ replaced by $B_{\rm pr}(f_{k},f_{l})$.
The PR prescription works equally well regardless of which of the two distributions is used for $R^{\alpha}$ [30]. The expression for the variance of biphase corresponding to phase-randomized bicoherence is similar to Eq. (4). It is estimated by using $b_{\rm pr}(f_{k},f_{l})$ instead of $b(f_{k},f_{l})$, and the variance of its biphase will be denoted as $\sigma_{\Phi}^{2}$.
@@
In this work, we focus on the auto--bicoherence and briefly present a feasiblity study on the cross--bicoherence in Appendix C. The main reason for this decision is that the cross--bicoherence analysis is computationally expensive when a priori it is unclear which detector systems’ perturbations would be phase-coupling to produce specific noise transients in the GW strain channel. Current detectors log these data in hundreds of thousands of auxiliary and physical-environmental-monitor channels. Even if one were to pick the top hundred suspects from that list for $x$ and $y$, computing the cross--bicoherence with $x\neq y$, and $z$ as the GW strain, involves ${\cal O}(10^{4})$ times more floating point operations than the auto--bicoherence, with all three data streams being the GW strain. Indeed, one can use certain well-motivated strategies, e.g., requiring the concurrence of transients in auxiliary channels with the GW strain data [33] to draw up such a shortlist. A second idea is to use a two-step hierarchical strategy, where in the first step one computes the auto-bicoherence on the GW strain channel at the locations of the noise transients. In the second step, one evaluates the more computationally expensive cross-bicoherence but only for that subset of transients from the first step that return the largest auto-bicoherence values. Combining both ideas will help reduce both the number of channels and the noise transients and can make the computation of the cross-bicoherence manageable. Details of that study will be presented in a future work [34]. For the rest of the paper, by bicoherence we will always mean the auto--bicoherence, unless otherwise noted.
Below we show with simulations how the bicoherence statistic can be used to discern the presence of quadratic phase coupling. Thereafter, we present some applications in real GW data. Throughout this paper, we fixed the values of some of the analysis parameters for both simulated signals and GW data, unless stated otherwise. The duration of all data chunks analyzed is chosen to be 4 sec, with a sampling frequency $f_{\rm s}=1024$ Hz; of course, the Nyquist frequency is set to $f_{N}=\frac{f_{\rm s}}{2}$. For computing the bicoherence, each time-series data chunk is divided into $M$ segments and their Fourier transforms taken after applying the Hanning window on them.
The optimal choice for the number of segments ($M$) is decided empirically. For a fixed chunk duration, if $M$ is small, then the relative contribution of noise to the bicoherence magnitude can be large. On the other hand, if $M$ is large, then the overlap of the segments will be large for the same chunk duration. In such a case, the different segments will cease to be independent and the bicoherence for a real signal, of a similar duration as the chunk, will plateau. By performing injection of simulated signals in real data and studying known short duration couplings (lasting no more than a few seconds) we found $M\approx 31$, which corresponds to a $98\%$ overlap among every pair of neighboring segments.
With the above specifications, for each 4s chunk the auto-bicoherence computation takes ${\cal O}(M{\cal N}^{3})$ floating-point operations (flop) with 8-byte long complex numbers, where ${\cal N}=4096$ frequency points. This adds up to about a tera-flop, which, along with read/write overheads, takes several minutes to execute on current processors. Thus, if the number of channels and noise transients can be kept manageable, the cross-bicoherence computation will be viable.